OpenStudy (anonymous):

The length of one leg of a right triangle is 20 ft and the length of the hypotenuse is 52 ft. What is the length of the other leg? A. 70.8 feet B. 5.7 feet C. 48.0 feet D. 32.0 feet

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

Use Pythagorean Theorem a2+b2=c2 202+b2=522 400+b2=2704 b2=2304 b=48

OpenStudy (jordanloveangel):

c

OpenStudy (anonymous):

Than you!

OpenStudy (jordanloveangel):

np